OpenStudy (shamil98):

\[\huge 4xy + 8x^2 + y^2 = 9x^3\] Differentiate implicitly. @Luigi0210

OpenStudy (kainui):

What's your best guess? I'll help you out! =D

OpenStudy (shamil98):

I know how to do it, not many questions atm, just having some fun :P

OpenStudy (kainui):

Oh, well I can do this with my eyes closed so I'll leave it alone haha.

OpenStudy (luigi0210):

Is it \[\LARGE 4y'+16x+2yy'=27x^2\]

OpenStudy (anonymous):

wow that fast i need help with one as well...i will post now

OpenStudy (shamil98):

Not quite, product rule for 4xy gives you 4y + 4xy'

OpenStudy (luigi0210):

Crap, forgot my math rules for a sec

OpenStudy (kc_kennylau):

\[\huge 4xy+8x^2+y^2=9x^3\]\[\huge4xy'+4y+16x+2yy'=27x^2\]\[\huge y'(4x+2y)=27x^2-4y-16x\]\[\huge y'=\frac{27x^2-4y-16x}{4x+2y}\]

OpenStudy (luigi0210):

;-;

OpenStudy (shamil98):

and we have a winner :)

OpenStudy (shamil98):

Good job @kc_kennylau

OpenStudy (kc_kennylau):

Merci :) @shamil98